OpenStudy (anonymous):

width of a rectangle is 3/4 of the lenght. perimeter of the rectangle becomes 50cm when the lenghr and width is increasedby 2cm find the width and lenght

myininaya (myininaya):

Let W be width Let L be length and if we need it let P be perimeter 2L+2W=P is formula to find perimeter W=3/4 * L P=50 when W+2 so we have have 2L+(3/4*L+2)=50

myininaya (myininaya):

oops mistake 2L+2(3/4*L+2)=50

OpenStudy (anonymous):

^what he said - I did area instead of perimeter.

OpenStudy (anonymous):

W = width, L = length. Perimeter =2W+2L W=3/4L. P=50 when W and L are incresed by 2. So... 2(3/4L+2)+2(L+2)=50 (3/2L+4)+(2L+4)=50 Combine like terms. 7/2L+8=50 Subtract 8 from both sides. 7/2L=42 Divide both sides by 7/2. L=42(2/7) L=12. W=3/4(12) W=9. Width =9cm and Length = 12cm.

myininaya (myininaya):

oh i didnt see that the length was also increased by 2 lol